Find the inverse of (4-x)/3
Answer by jim_thompson5910(28595)

y = (4-x)/3


x = (4-y)/3


3x = 4-y


3x - 4 = -y


-3x + 4 = y


y = -3x + 4


So the inverse is -3x + 4


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim